OpenStudy (anonymous):

Simplify the expression 5^3 × 5^8.

OpenStudy (igreen):

We add the exponents when we're multiplying like bases.

OpenStudy (igreen):

\(\sf 5^3 \times 5^8 \rightarrow 5^{3+8}\rightarrow~ \! ?\)

OpenStudy (anonymous):

ohh okay thanks for the help @iGreen

OpenStudy (igreen):

Okay, what did you come up with?

OpenStudy (anonymous):

\[5^{11}\] @iGreen

OpenStudy (unklerhaukus):

this works because \[5^3\times5^8\\=(5\times5\times5)\times (5\times5\times5\times5\times5\times5\times5\times5)\\ =5\times5\times5\times5\times5\times5\times5\times5\times5\times5\times5\\=5^?\]

OpenStudy (igreen):

Yes, that is correct.

OpenStudy (igreen):

\(\color{lime}{\underbrace{\overbrace{\Huge\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\\\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\\\Large{~\succ}\bigstar\boxed{\sf Welcome~to~\color{#3dabea}{Open}\color{#9cc575}{Study}!!!}\bigstar\hspace-5px\prec\\\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\\\Huge{\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast}}}}\) \(\small\sf\color{red}{Made~by~@nobrainiachere,~edited~by~@iGreen}\) You can give medals by clicking 'Best Response'.